Schell Company manufactures automobile floor mats. It currently has two product lines, the Standard and the Deluxe. Suppose that Schell has conducted further research into its overhead and potential cost drivers. As a result, the company has compiled the following detailed information, breaking total overhead into three cost pools:

Activity Cost Pools Cost Driver Cost
Assigned to Pool
Quantity/Amount
Consumed by
Standard Floor Mat Line
Quantity/Amount
Consumed by
Deluxe Floor Mat Line
Material handling Number of moves $ 3,094.00 44 moves 47 moves
Quality control Number of inspections $ 7,688.00 560 inspections 680 inspections
Machine maintenance Number of machine hours $ 13,718.00 3,040 machine hours 4,180 machine hours


Required:

1. Calculate the activity rates for each cost pool assuming Schell uses an ABC system.

2. Calculate the amount of overhead that Schell will assign to the Standard floor mat line.

3. Determine the amount of overhead Schell will assign to the Deluxe product line.

Required information

Bunker makes two types of briefcase, fabric and leather. The company is currently using a traditional costing system with labor hours as the cost driver but is considering switching to an activity-based costing system. In preparation for the possible switch, Bunker has identified two activity cost pools: materials handling and setup. Pertinent data follow:

Fabric Case Leather Case
Number of labor hours 14,000 9,000
Number of material moves 540 810
Number of setups 48 72

Total estimated overhead costs are $203,550, of which $168,750 is assigned to the materials handling cost pool and $34,800 is assigned to the setup cost pool.

Required:
1.
Calculate the overhead assigned to the leather case line using the traditional costing system based on direct labor hours.
2. Calculate the overhead assigned to the leather case line using ABC.
3. Was the leather case over- or undercosted by the traditional cost system compared to ABC?
Answer #1

1) Calculate activity rate

Activity rate
Material handling 3094/91 = 34 per move
Quality control 7688/1240 = 6.20 per inspection
Machine maintenance 13718/7220 = 1.90 per machine hour

1a) Overhead assigned to Standard floor mat line

Overhead assigned
Material handling 1496
Quality control 3472
Machine maintenance 5776
Total 10744

1b) Overhead assigned to Deluxe floor mat line

Overhead assigned
Material handling 1598
Quality control 4216
Machine maintenance 7942
Total 13756

2a) Overhead assigned to Fabric under traditional costing = 203550/23000*9000 = 79650

2b) Overhead assigned to Fabric under ABC = (810*125+290*72) = 122130

2c) Under costed under Traditional costing by 42480